OpenStudy (anonymous):

solve for the variable . -x/9=5/3

jimthompson5910 (jim_thompson5910):

You want to isolate x, so you want to move that 9 over. So you need to undo dividing by 9 and do what to both sides?

OpenStudy (anonymous):

uhm not sure

jimthompson5910 (jim_thompson5910):

What is the opposite of division?

OpenStudy (anonymous):

multiply both sides by 9

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so -x/9=5/3 (-x/9)*9=(5/3)*9 -x = (5/3)*(9/1) -x = (5*9)/(3*1) -x = 45/3 -x = 15 So because -x = 15, this means x = -15 (just flip the signs of each)